Jefferson Lab HALL A Logbook for January 01 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 56209 01/31/01 23:03 Lingyan left-arm HV trip
  • 56196 01/31/01 15:26 atrig Trigger-Download
  • 56193 01/31/01 15:14 atrig Trigger-Download
  • 56192 01/31/01 15:00 Lingyan left arm HV restored
  • 56191 01/31/01 14:58 R. Michaels roc2 power reboot
  • 56186 01/31/01 12:32 H. Xiang HV's for L-Arm Restored
  • Tuesday

  • 56183 01/30/01 18:42 bogdanw Left Arm HV
  • 56181 01/30/01 00:02 W. Boeglin Shift summary
  • Monday

  • 56174 01/29/01 22:41 Adaq Angle Setting for kinematics 10
  • 56167 01/29/01 20:37 ddutta Kine-10 kaons
  • 56164 01/29/01 19:24 markowitz kin 4 kaons
  • 56161 01/29/01 19:14 markowitz kin 6 run 1852 kaons
  • 56160 01/29/01 19:11 markowitz kaons kin 6 run 1850
  • 56159 01/29/01 19:09 markowitz kaons kin 6 run 1849
  • 56148 01/29/01 16:01 LeRose Shift Summary
  • 56146 01/29/01 15:52 jpchen missing mass plot for KIN 8 run 1878
  • 56145 01/29/01 15:14 LeRose Left VDC Discriminator Levels Changed
  • 56142 01/29/01 14:48 jpchen KIN 8 missing mass plot for RUN 1878
  • 56141 01/29/01 14:43 jpchen KIN 9 missing mass plot Run 1869
  • 56140 01/29/01 13:50 jpchen Left VDC multiplicity variation
  • 56127 01/29/01 12:24 markowitz optics run 1864 kin 9
  • 56126 01/29/01 12:17 jpchen missing mass plot for KIN 7 run 1858
  • 56125 01/29/01 12:12 jpchen KIN 7 run 1857 massing mass plot
  • 56122 01/29/01 11:52 markowitz optics target analysis
  • 56119 01/29/01 11:48 markowitz kin 7 missing mass plot
  • 56110 01/29/01 11:03 markowitz spot at kinematics 9
  • 56104 01/29/01 10:39 LeRose Kinematics 9 Magnets
  • 56099 01/29/01 09:53 LeRose BeO
  • 56095 01/29/01 09:22 jpchen KIN7 run 1857 missing mass plot
  • 56094 01/29/01 09:16 jpchen more KIN 6 missing mass plot
  • 56093 01/29/01 09:12 LeRose Left HV Trip
  • 56091 01/29/01 09:10 jpchen KIN 6 missing mass plots
  • 56087 01/29/01 08:21 LeRose Day Shift
  • 56086 01/29/01 08:00 reinhold shift summary
  • 56084 01/29/01 07:38 dohrmann Detector and physics replay remarks
  • 56080 01/29/01 07:27 Deur VDC wire efficiency
  • 56082 01/29/01 07:25 reinhold MCC: box supply trip
  • 56079 01/29/01 07:01 reinhold HV trip run 1853
  • 56075 01/29/01 06:45 dohrmann kin6, run 1850 kaons
  • 56074 01/29/01 06:43 reinhold run 1854: pedestal run
  • 56073 01/29/01 06:33 reinhold MCC does diagnostics
  • 56071 01/29/01 06:31 reinhold run 1855: move to kin 7
  • 56069 01/29/01 06:28 adaq kaons run 1848 , kin6
  • 56068 01/29/01 06:23 dohrmann missingmass_kin6.kumac
  • 56060 01/29/01 04:32 dohrmann kaon total kin 5
  • 56058 01/29/01 04:17 reinhold run 1849: lower current back to 80uA
  • 56054 01/29/01 03:51 reinhold try to increase beam current
  • 56053 01/29/01 03:48 adaq run 1839 kaons
  • 56051 01/29/01 03:37 reinhold move to 15 cm LH2
  • 56049 01/29/01 03:30 adaq run 1838 kaons
  • 56039 01/29/01 03:05 reinhold moved to dummy 15 cm target
  • 56037 01/29/01 03:00 reinhold run 1841 optics for KIN 6
  • 56033 01/29/01 02:53 reinhold run 1840 optics KIN 5
  • 56032 01/29/01 02:52 adaq run 1837 kaons
  • 56029 01/29/01 02:30 J.Lenoble high current in some shower elements
  • 56028 01/29/01 02:19 adaq run 1836 junk !!
  • 56026 01/29/01 02:14 adaq run 1835 kaons
  • 56025 01/29/01 02:11 Adaq run 1836 junk ??
  • 56022 01/29/01 01:14 adaq run 1832
  • 56019 01/29/01 00:52 adaq run 1834
  • 56017 01/29/01 00:08 G. M. Urciuo Swing Shift Report
  • Sunday

  • 55984 01/28/01 16:15 comanm right,left y target
  • 55981 01/28/01 16:03 Kanda Shift Summary
  • 55973 01/28/01 15:14 markowitz kin 5, beo, 9-foil, dummy runs done
  • 55971 01/28/01 15:12 markowitz bogies_LEFT
  • 55969 01/28/01 15:04 baker cryotarget ioc reboot
  • 55965 01/28/01 14:44 markowitz BeO
  • 55964 01/28/01 14:37 markowitz magnet screen, setting kin 5
  • 55962 01/28/01 13:15 comanm hv right arm
  • 55959 01/28/01 12:54 Kanda HV for AC tripped
  • 55957 01/28/01 11:51 Kanda Sudden Trigger rate decrease
  • 55949 01/28/01 08:48 markowitz spot screenshot
  • 55946 01/28/01 07:59 dohrmann Shift Summary
  • 55944 01/28/01 07:43 dohrmann VDC eff. and correlation
  • 55943 01/28/01 07:25 Deur Scintillator correlations
  • 55939 01/28/01 07:13 adaq left spec vdc efficiency
  • 55935 01/28/01 06:29 pking Kin4 kaons through run 1804
  • 55933 01/28/01 05:32 dohrmann ps8
  • 55932 01/28/01 05:24 dohrmann HV for left hrs s2.adc_r_c2 plugged in
  • 55913 01/28/01 02:49 Jack Crate resets.
  • 55911 01/28/01 02:06 Jack Left arm slow controls.
  • 55910 01/28/01 01:39 dohrmann status update slow control problem left arm
  • 55909 01/28/01 01:11 dohrmann Total number of kaons for 1771,1772,1774
  • 55908 01/28/01 00:45 comanm missing mass kinematic 4
  • 55906 01/28/01 00:04 G. M . Urciu C shift report
  • Saturday

  • 55905 01/27/01 23:36 R. Michaels more DAQ woes... answer
  • 55896 01/27/01 21:58 R. Michaels roc 2 problems
  • 55883 01/27/01 20:52 R. Michaels Problems TS1 crate, fixed
  • 55881 01/27/01 20:15 markowitz hv crate failure and widening tc gate
  • 55878 01/27/01 19:54 R. Michaels increased coinc width, portserver ok
  • 55875 01/27/01 19:39 atrig Trigger-Download
  • 55873 01/27/01 18:35 atrig Trigger-Download
  • 55872 01/27/01 17:48 markowitz HV trip realization
  • 55871 01/27/01 17:44 kramer left high voltage tripped
  • 55868 01/27/01 17:27 R. Michaels L-arm portserver
  • 55867 01/27/01 17:14 Adaq hatsv4
  • 55865 01/27/01 16:14 C. C. Chang Run Summary
  • 55859 01/27/01 12:38 Kanda Missing mass for Kinematics #3 (revised)
  • 55857 01/27/01 12:34 markowitz run 1771 -- first good kin 4
  • 55855 01/27/01 12:33 markowitz starting run 1770 (kin 4)
  • 55853 01/27/01 12:27 markowitz moving to hydrogen
  • 55851 01/27/01 12:18 markowitz mcc cleanup results
  • 55849 01/27/01 12:07 markowitz putting in 15 cm dummy
  • 55847 01/27/01 11:56 Adaq magnet screen with kinematic 4 settings
  • 55845 01/27/01 11:37 markowitz setting up for kin 4
  • 55844 01/27/01 11:35 Kanda Missing mass for Kinematics #3
  • 55840 01/27/01 09:39 reinhold wrong fig in 55839
  • 55839 01/27/01 09:33 reinhold radiative corrections for KIN 3
  • 55838 01/27/01 09:31 C. C. Chang Acceptances
  • 55837 01/27/01 09:16 Chang Run Summary
  • 55836 01/27/01 08:53 reinhold kinematic 1 statistics
  • 55833 01/27/01 08:17 pking Run 1765 scintillator TDC spectra
  • 55832 01/27/01 08:05 reinhold shift summry
  • 55830 01/27/01 07:31 reinhold KIN 3 statistics and cross section
  • 55828 01/27/01 06:52 reinhold run 1764 DAQ crash
  • 55821 01/27/01 00:59 reinhold run1760
  • 55820 01/27/01 00:56 reinhold run 1759
  • 55815 01/27/01 00:21 reinhold run 1749 wrong header
  • 55814 01/27/01 00:06 reinhold run 1758 wrong header
  • 55811 01/27/01 00:04 Raue Change of angles
  • Friday

  • 55808 01/26/01 22:33 markowitz prescales set for optics
  • 55807 01/26/01 22:10 Kramer CODA died
  • 55805 01/26/01 21:10 Adaq magnet settings
  • 55803 01/26/01 19:57 Raue data problem
  • 55800 01/26/01 19:48 Raue data problem follow up
  • 55799 01/26/01 19:48 K. McCormick New ESR variable added
  • 55798 01/26/01 19:41 Raue beam stability
  • 55797 01/26/01 19:17 Raue Data problem
  • 55791 01/26/01 16:05 Saha Shift Summary
  • 55787 01/26/01 14:37 gilman BeO check
  • 55788 01/26/01 14:34 Adaq ytarget for kinemtics 1
  • 55785 01/26/01 13:33 atrig Trigger-Download
  • 55783 01/26/01 10:14 reinhold missing mass background subtracted
  • 55782 01/26/01 09:41 gilman start of run 1750 wrong
  • 55779 01/26/01 09:14 R. Michaels obscure note about T4
  • 55778 01/26/01 08:23 reinhold missing mass how-to
  • 55777 01/26/01 08:07 reinhold shift summary
  • 55775 01/26/01 07:38 Markowitz kaon rates
  • 55773 01/26/01 07:19 Adaq that missing mass spectrum
  • 55771 01/26/01 06:55 reinhold beam for H2 production
  • 55769 01/26/01 06:46 reinhold moved target to 15 cm H2
  • 55768 01/26/01 06:46 reinhold LEFT HRS to KIN 2
  • 55767 01/26/01 06:43 reinhold set RIGHT HRS to KIN 2
  • 55763 01/26/01 05:59 reinhold first lambda missing mass
  • 55760 01/26/01 05:14 reinhold run 1744 is junk
  • 55759 01/26/01 05:14 reinhold prescale for optics
  • 55755 01/26/01 05:06 reinhold beam for optics target
  • 55752 01/26/01 04:57 reinhold moved to 9 foil optics target
  • 55749 01/26/01 04:06 reinhold moved to 15cm dummy target
  • 55747 01/26/01 03:08 pking HV Alarm on LARM pion rejector
  • 55744 01/26/01 02:08 pking LARM buttons crash "pmthv"
  • 55741 01/26/01 01:06 pking Left arm Aero A2 ADC25 Pedestal
  • 55739 01/26/01 01:03 Raue Shift Summary
  • Thursday

  • 55712 01/25/01 16:03 A. Saha Shift Summary
  • 55710 01/25/01 15:08 Saha Beam Spot on BeO target - picture
  • 55707 01/25/01 14:32 Saha Beam back, beam spot on BeO target
  • 55704 01/25/01 12:59 krishniw Beta in the right arm
  • 55703 01/25/01 12:53 krishniw Beta in the left arm
  • 55701 01/25/01 12:21 R. Michaels Mods to trigger/DAQ
  • 55700 01/25/01 12:13 Saha New BCM Charge Coefficients
  • 55698 01/25/01 11:32 atrig Trigger-Download
  • 55696 01/25/01 11:30 Reinhold spectrometer setting screen shot
  • 55693 01/25/01 11:25 Reinhold magnet setting
  • 55689 01/25/01 11:15 Beise target
  • 55688 01/25/01 11:13 Markowitz adaq root partition full/ espace errors
  • 55687 01/25/01 11:08 atrig Trigger-Download
  • 55686 01/25/01 10:57 Reinhold set spectrometer angles
  • 55685 01/25/01 10:55 atrig Trigger-Download
  • 55682 01/25/01 10:04 Markowitz Pion rejector HV turned off
  • 55674 01/25/01 09:06 Saha Beam Spot on BeO target
  • 55672 01/25/01 08:59 Saha Beam Spot on BeO Target
  • 55671 01/25/01 08:58 Adev Beam Spot
  • 55667 01/25/01 08:02 G. M. Urciuo Hall A owl shift summary
  • 55651 01/25/01 05:38 atrig Trigger-Download
  • 55627 01/25/01 02:14 Lingyan Scintillator TDC after width change(right)
  • 55626 01/25/01 02:07 Lingyan Scintilaltor TDC spectra after width change(left)
  • 55620 01/25/01 00:47 Lingyan keep regulator on
  • Wednesday

  • 55616 01/24/01 23:56 S. Sirca Evening shift summary
  • 55615 01/24/01 23:50 Higinbotham EP Data
  • 55612 01/24/01 23:09 Lingyan on LD2
  • 55611 01/24/01 23:04 Higinbotham Magnet Screen
  • 55609 01/24/01 22:52 Higinbotham Gas Cherenkov Voltage Change
  • 55607 01/24/01 22:49 Higinbotham Harp Scan
  • 55603 01/24/01 21:42 S. Sirca setspec ambiguity
  • 55602 01/24/01 21:35 S. Sirca Loop2 fan alarms
  • 55599 01/24/01 20:40 S. Sirca Feature: crash on resize
  • 55598 01/24/01 20:38 S. Sirca Trigger download for coinc kin #17
  • 55597 01/24/01 20:31 atrig Trigger-Download
  • 55596 01/24/01 20:30 atrig Trigger-Download
  • 55595 01/24/01 19:19 Jack CAMAC and Fastbus filters
  • 55594 01/24/01 19:18 Jack S2R4 PMT
  • 55593 01/24/01 18:51 R. Michaels re: double peaks in TDCs
  • 55582 01/24/01 16:28 Adaq summary of polarities up to run # 1677
  • 55579 01/24/01 15:40 weinstein day shift summary
  • 55570 01/24/01 14:20 X. Jiang left-arm HV screen during run 1677
  • 55556 01/24/01 11:14 K. McCormick Cryotarget cryostat screen
  • 55553 01/24/01 10:40 Beise LD2 Target fan speed was low
  • 55549 01/24/01 10:31 Lingyan unstable L-gas #9
  • 55545 01/24/01 08:50 Beise Target Controls GUI
  • 55543 01/24/01 08:25 Adaq Noisy channel on left-arm gas
  • 55542 01/24/01 08:17 weinstein reboot target ioc
  • 55539 01/24/01 08:05 Dutta Shift summary
  • 55538 01/24/01 07:36 ddutta VDC Disc level was low
  • 55528 01/24/01 06:16 atrig Trigger-Download
  • 55525 01/24/01 06:02 ddutta spectrometer moved
  • 55517 01/24/01 02:29 ddutta Taarget IOC reboot
  • 55509 01/24/01 00:13 S. Sirca Evening shift summary
  • 55505 01/24/01 00:05 S. Sirca HV alarm handler malfunction?
  • Tuesday

  • 55501 01/23/01 23:14 S. Sirca ROC1 crash and consequences
  • 55478 01/23/01 16:44 S. Sirca Left-HRS magnets ready
  • 55477 01/23/01 15:57 weinstein day shift summary
  • 55475 01/23/01 15:03 weinstein left q2 quench again
  • 55468 01/23/01 13:52 weinstein q2 left quench again
  • 55467 01/23/01 13:48 Saha Hall A Arc Energy measurement
  • 55466 01/23/01 13:35 weinstein q2 left quench
  • 55465 01/23/01 10:12 weinstein dipole liquid level alarm
  • 55463 01/23/01 08:14 Lingyan shift summary
  • 55461 01/23/01 07:39 Lingyan post-ESR crash status
  • 55459 01/23/01 00:01 S. Sirca Evening shift summary
  • 55456 01/23/01 00:00 S. Sirca Cryo targets status
  • Monday

  • 55455 01/22/01 21:18 S. Sirca Post-ESR crash status
  • 55453 01/22/01 20:25 ddutta ESR crash
  • 55452 01/22/01 19:37 S. Sirca Hall C requests half-wave plate IN
  • 55446 01/22/01 17:57 Lingyan more on zero-vdc events
  • 55444 01/22/01 17:29 S. Sirca Left-HRS Q2 goes off
  • 55440 01/22/01 16:12 weinstein shift summary
  • 55435 01/22/01 14:38 x. jiang right-arm x vs. y distribution at VDC
  • 55433 01/22/01 13:37 x.jiang VDC multiplicity in left arm
  • 55432 01/22/01 13:34 jpchen cryotarget temperature/pressure oscillation
  • 55428 01/22/01 11:44 X. Jiang zero_VDC hits events in gammap99 data
  • 55427 01/22/01 11:41 X. Jiang adjust high voltage in pion rejector
  • 55424 01/22/01 11:24 krishniw Left arm S2 right 1,2,3
  • 55423 01/22/01 10:48 jpchen Re-started target GUI
  • 55418 01/22/01 09:53 krishniw After changing PMT's on S2 right, 1,2 ,4
  • 55415 01/22/01 08:24 Lingyan Zhu shift summary
  • 55397 01/22/01 01:35 X. Jiang Adjusted high voltage for left-arm pion rejector.
  • 55392 01/22/01 00:13 S. Sirca Evening shift summary
  • Sunday

  • 55385 01/21/01 23:01 atrig Trigger-Download
  • 55369 01/21/01 20:52 bogdanw S1/S2 PMT checkout
  • 55370 01/21/01 20:48 Strauch Left-HRS Q2 magnet went off
  • 55368 01/21/01 20:15 ddutta VDC zero hits - the saga continues
  • 55365 01/21/01 19:46 Strauch Satus screen for coincidence kinematic 16
  • 55363 01/21/01 19:04 x. jiang pion rejector amplifier channel change.
  • 55362 01/21/01 17:51 ddutta VDC-hits events with radiator out run
  • 55361 01/21/01 16:24 Feng Shift summary
  • 55359 01/21/01 15:39 x. jiang zero VDC-hits events with radiator out run.
  • 55358 01/21/01 12:36 K. McCormick Beamline HV crate switched out
  • 55357 01/21/01 11:55 ddutta Status Update
  • 55356 01/21/01 09:13 R. Michaels dont remove /tmp/et_sys_tst1
  • 55354 01/21/01 08:16 Lingyan shift summary
  • 55352 01/21/01 06:39 Lingyan beam off
  • 55336 01/21/01 00:17 S. Sirca Evening shift summary
  • 55334 01/21/01 00:14 S. Sirca HOWTO get out of a CODA crash
  • Saturday

  • 55332 01/20/01 23:43 R. Michaels re: starting CODA after crashes
  • 55322 01/20/01 20:15 ddutta loop 2 JT valve
  • 55304 01/20/01 16:27 H. Gao Shift summary
  • 55267 01/20/01 11:49 atrig Trigger-Download
  • 55264 01/20/01 11:43 seely target IOC reboot
  • 55259 01/20/01 11:05 ddutta resetting the Beamline HV Crate
  • 55247 01/20/01 08:36 Lingyan shift summary
  • 55245 01/20/01 08:17 seely replay on adaql1
  • 55242 01/20/01 08:05 Lingyan about online replay
  • 55220 01/20/01 05:11 comanm coda hangup
  • 55213 01/20/01 01:18 Lingyan l-arm Q2 nitrogen back
  • 55206 01/20/01 00:12 S. Sirca Evening shift summary
  • 55207 01/20/01 00:09 S. Sirca Left HRS Q2 nitrogen low
  • Friday

  • 55204 01/19/01 22:53 S. Sirca Raster IOC crash
  • 55201 01/19/01 22:15 S. Sirca Begin kinematics #17
  • 55194 01/19/01 19:52 S. Sirca Starting with run # 1526, radiator OUT
  • 55187 01/19/01 18:40 S. Sirca Starting with run # 1523, radiator IN
  • 55184 01/19/01 17:36 X. Jiang noticed zero hits events on VDC
  • 55178 01/19/01 16:41 S. Sirca Accelerator re-tune
  • 55176 01/19/01 16:38 S. Sirca Starting with run # 1519, radiator is OUT.
  • 55168 01/19/01 15:38 gilman approved for 50 microamps
  • 55153 01/19/01 13:36 gilman coda crash on roc 14 run 1508
  • 55149 01/19/01 12:57 R. Holt target IOC reboot
  • 55144 01/19/01 12:46 gilman status and run sheets
  • 55135 01/19/01 11:30 ddutta PHP script
  • 55128 01/19/01 08:23 Lingyan shift summary
  • 55126 01/19/01 08:07 Lingyan HV monitoring script
  • 55120 01/19/01 04:46 Lingyan left arm HV problem
  • 55111 01/19/01 02:52 Adaq He level oscillation
  • 55092 01/19/01 00:10 S. Sirca Evening shift summary
  • 55094 01/19/01 00:04 seely LD2 fan frequency
  • Thursday

  • 55073 01/18/01 21:32 adaq re:Hit-pattern-T1
  • 55070 01/18/01 21:15 S. Sirca Raster problem
  • 55063 01/18/01 20:02 Lingyan xy cut on scintillator adc
  • 55060 01/18/01 19:54 Lingyan xy cut on scintillator adc
  • 55059 01/18/01 19:53 S. Sirca Radiator IN starting with run # 1470
  • 55057 01/18/01 19:49 S. Sirca Coolant problems persist
  • 55049 01/18/01 18:04 S. Sirca Radiator OUT starting with run # 1466
  • 55048 01/18/01 17:52 S. Sirca Left D1 coolant level
  • 55043 01/18/01 17:10 S. Sirca Left D1 coolant level fluctuations
  • 55042 01/18/01 17:05 Lingyan Hit pattern for T1 or T5 on R-arm
  • 55033 01/18/01 16:18 LeRose Shift summary
  • 55025 01/18/01 15:13 LeRose Beam back
  • 55023 01/18/01 15:07 LeRose RF problem
  • 55010 01/18/01 13:53 C. Williamso cerenkov mirror
  • 55008 01/18/01 13:51 C. Williamso cerenkov a1 a2
  • 55011 01/18/01 13:43 K. McCormick HV problems
  • 55006 01/18/01 13:18 LeRose Raster Problem
  • 55005 01/18/01 13:14 LeRose Energy Measurement Aborted
  • 55003 01/18/01 12:02 LeRose Energy Measurement
  • 55001 01/18/01 11:46 LeRose Beam Back
  • 54999 01/18/01 11:19 LeRose MCC taking beam
  • 54995 01/18/01 11:02 Saha Arc-Integral Measurement
  • 54991 01/18/01 10:42 LeRose Beam Back
  • 54990 01/18/01 10:20 LeRose Beam Off
  • 54985 01/18/01 10:12 R. Holt ADC spectra
  • 54976 01/18/01 08:53 LeRose Right Collimator
  • 54969 01/18/01 08:14 mitchell shift summary
  • 54956 01/18/01 06:22 Lingyan Scintillator trigger inefficiency and contamination
  • 54946 01/18/01 04:02 atrig Trigger-Download
  • 54943 01/18/01 03:44 atrig Trigger-Download
  • 54934 01/18/01 00:39 Lingyan dirty left gas adc
  • 54932 01/18/01 00:10 E.Chudakov Shift Summary
  • Wednesday

  • 54929 01/17/01 23:15 Lingyan Scintillator threshold(left arm)
  • 54928 01/17/01 23:12 Lingyan Scintillator threshold(right arm)
  • 54916 01/17/01 17:04 R. Michaels day shift summary
  • 54914 01/17/01 16:03 R. Michaels re-download trigger
  • 54912 01/17/01 15:57 atrig Trigger-Download
  • 54910 01/17/01 15:47 krishniw Corelations of Preshower vs Shower
  • 54909 01/17/01 15:43 krishniw Shower Preshower and Gas cerenkov
  • 54907 01/17/01 14:36 krishniw target endcaps can be seen
  • 54905 01/17/01 13:41 Adaq Target IOC reboot
  • 54901 01/17/01 13:17 krishniw preshower working at 5.6 GeV
  • 54900 01/17/01 13:12 krishniw Preshower working at 5.6 GeV
  • 54896 01/17/01 12:04 R. Michaels T5, to correct for pulser
  • 54893 01/17/01 10:58 R. Michaels DAQ synch check
  • 54891 01/17/01 08:38 R. Michaels trigger observations
  • 54889 01/17/01 08:33 ddutta CO2 pressure in gas shed
  • 54887 01/17/01 08:16 mitchell shift summary
  • 54879 01/17/01 05:03 kramerk target IOC reboot
  • 54876 01/17/01 04:07 kramerk CO2 gas shed pressure -74psi
  • 54873 01/17/01 03:21 Lingyan radiator linearity check
  • 54862 01/17/01 00:06 Adaq Shift Summary
  • Tuesday

  • 54846 01/16/01 20:04 Feng LH2 Fan Frequency Can Be Set At 25%(60 Hz)
  • 54841 01/16/01 19:47 Feng LD2 Fan Frequency Should NOT Be Set Above 20%(48 Hz)
  • 54839 01/16/01 19:43 E.Chudakov Starting up
  • 54834 01/16/01 18:47 wangxu Trigger check(2)
  • 54831 01/16/01 18:44 Adaq Trigger check(1)
  • 54826 01/16/01 18:24 wang xu Scaler software and bcm check
  • 54821 01/16/01 16:57 atrig Trigger-Download
  • 54820 01/16/01 16:08 krishniw A2 proton detection
  • 54817 01/16/01 15:48 atrig Trigger-Download
  • 54816 01/16/01 15:39 atrig Trigger-Download
  • 54815 01/16/01 15:25 x. jiang be careful with high voltage control !!!
  • 54814 01/16/01 15:21 Adaq magnet display screen
  • 54813 01/16/01 15:19 R. Michaels trigger issues (please read)
  • 54811 01/16/01 14:29 krishniw Particle id from S1/S2
  • 54810 01/16/01 14:25 C. Williamso Hall A magnets
  • 54809 01/16/01 14:17 krishniw delta
  • 54808 01/16/01 13:50 jpchen Hall A target cool-down completed
  • 54804 01/16/01 07:48 kramerk Shift Summary
  • 54803 01/16/01 00:08 R. Michaels retiming defined by S2-R. And T0 change
  • Monday

  • 54786 01/15/01 22:48 atrig Trigger-Download
  • 54785 01/15/01 21:29 E.Chudakov Recovery status
  • 54783 01/15/01 18:56 kramerk ff
  • 54781 01/15/01 18:54 R. Michaels trigger checks (after bug fix)
  • 54782 01/15/01 18:53 Jack PMT #9 base on left arm gas c
  • 54780 01/15/01 18:45 R. Michaels fastbus busy, synch, fb lib errors
  • 54777 01/15/01 17:58 atrig Trigger-Download
  • 54774 01/15/01 17:56 atrig Trigger-Download
  • 54771 01/15/01 17:53 atrig Trigger-Download
  • 54769 01/15/01 17:51 atrig Trigger-Download
  • 54765 01/15/01 17:22 bogdanw HV controls confusion
  • 54764 01/15/01 16:27 R. Michaels trigger problems
  • 54761 01/15/01 15:59 Meekins New lifter vernier readings.
  • 54758 01/15/01 15:55 Lingyan dhist and dataspy
  • 54757 01/15/01 15:49 x. jiang adjustment of pion rejector hv
  • 54750 01/15/01 15:27 atrig Trigger-Download
  • 54748 01/15/01 14:36 Lingyan start-run entry
  • 54746 01/15/01 14:17 Saha BCM coefficients
  • 54745 01/15/01 11:43 atrig Trigger-Download
  • 54744 01/15/01 11:16 jpchen ESR recovering, magnets and targets cool-down
  • 54741 01/15/01 06:48 kramerk Request for leak check
  • 54740 01/15/01 06:33 dohrmann shift summary
  • 54739 01/15/01 05:57 dohrmann targets and magnets
  • 54738 01/15/01 05:45 dohrmann ESR down
  • 54737 01/15/01 05:15 dohrmann ESR coolant apparently gone
  • 54732 01/15/01 00:18 dohrmann Trigger #17 downloaded
  • 54731 01/15/01 00:17 atrig Trigger-Download
  • 54729 01/15/01 00:06 bogdanw shift summary
  • Sunday

  • 54728 01/14/01 21:25 seely preshower on cosmic
  • 54726 01/14/01 20:41 bogdanw Short GC PMT 9
  • 54725 01/14/01 19:29 seely control screen
  • 54724 01/14/01 19:28 seely control screen
  • 54723 01/14/01 18:31 X. Zheng gas cherenkov, left arm, mirror #9 status
  • 54721 01/14/01 17:10 X. Zheng Gas cherenkov, left arm, mirror #9 problem
  • 54719 01/14/01 16:14 Lingyan Shift Summary
  • 54717 01/14/01 13:27 Lingyan a2 adcsum
  • 54716 01/14/01 13:23 Lingyan aerogel calibration & a1, am adcsum
  • 54713 01/14/01 11:50 atrig Trigger-Download
  • 54711 01/14/01 11:42 atrig Trigger-Download
  • 54697 01/14/01 08:13 Hong Xiang Shift summary
  • 54672 01/14/01 03:42 dohrmann ROC 2 interrupt message
  • 54671 01/14/01 03:32 adaq reboot ROC2 restart CODA
  • 54668 01/14/01 03:16 adaq components window ROC2/hallasfi2
  • 54665 01/14/01 02:30 adaq changed to kinematics 2
  • 54663 01/14/01 01:47 x. jiang adjusted hv settings on pion-rejector
  • 54662 01/14/01 01:44 dohrmann TS changed HV of 4 chan
  • 54657 01/14/01 00:40 Lingyan start-run entry
  • 54653 01/14/01 00:18 bogdanw shift summary
  • Saturday

  • 54642 01/13/01 23:08 seely control screen
  • 54641 01/13/01 23:08 seely control screen
  • 54632 01/13/01 20:33 R. Michaels re: T5 double peaks
  • 54629 01/13/01 19:55 R. Michaels probable origin of weird TDC timing
  • 54622 01/13/01 17:46 x. jiang changed adc cable connnection on pion-rejector layer1-block-22.
  • 54618 01/13/01 16:56 seely control screen
  • 54610 01/13/01 16:25 Calarco Shift summary
  • 54601 01/13/01 15:31 Meekins Target Lifter Failure
  • 54599 01/13/01 15:14 Lingyan detmap change of PR
  • 54600 01/13/01 15:12 Calarco magnet settings
  • 54598 01/13/01 14:50 atrig Trigger-Download
  • 54596 01/13/01 13:10 R. Michaels DAQ synch check software
  • 54595 01/13/01 12:49 Calarco magnet settings
  • 54586 01/13/01 11:56 atrig Trigger-Download
  • 54585 01/13/01 11:09 Calarco Target problem
  • 54584 01/13/01 10:42 Calarco magnet settings
  • 54569 01/13/01 08:12 kinney A-shift Summary
  • 54557 01/13/01 03:34 atrig Trigger-Download
  • 54543 01/13/01 00:21 bogdanw shift summary
  • 54542 01/13/01 00:20 R. Michaels reminder: trigger
  • 54545 01/13/01 00:12 Lingyan HV change for right S1
  • 54540 01/13/01 00:05 atrig Trigger-Download
  • Friday

  • 54534 01/12/01 23:22 X. Zheng Gas cherenkov mirrors, right arm
  • 54533 01/12/01 23:18 X. Zheng Gas cherenkov mirrors, left arm
  • 54532 01/12/01 23:11 X. Zheng Database changed for gas cherenkov
  • 54527 01/12/01 21:46 X.Zheng Gas cherenkov performance
  • 54514 01/12/01 19:16 Adaq hv
  • 54510 01/12/01 18:55 gilman radiator in runstart, fastbus error
  • 54508 01/12/01 18:41 seely control screen
  • 54505 01/12/01 18:31 mitchell ep energy measurement
  • 54501 01/12/01 17:32 Nanda BCM caibration
  • 54499 01/12/01 16:43 Calarco ROC2 interrupts
  • 54497 01/12/01 16:30 R. Michaels various problems
  • 54496 01/12/01 16:27 gomez bcm/reboot hallasc5
  • 54489 01/12/01 15:41 Saha BPM/Harp Calibrations
  • 54487 01/12/01 15:31 Saha Arc Energy Measurement
  • 54456 01/12/01 13:52 R. Michaels L and R arm scaler config
  • 54455 01/12/01 13:50 adaq right-arm-scaler-config
  • 54454 01/12/01 13:50 adaq left-arm-scaler-config
  • 54453 01/12/01 13:29 Saha BCM Calibrations
  • 54439 01/12/01 12:04 John A Shower and Preshowere calibrations
  • 54436 01/12/01 11:32 Gomez re: hallasc16 & dipole prob.
  • 54430 01/12/01 10:50 Saha Arc-Integral measurement
  • 54429 01/12/01 10:43 Saha Harp Scans
  • 54422 01/12/01 08:56 Meekins New vernier positions measured 1/12/01
  • 54421 01/12/01 08:05 atrig Trigger-Download
  • 54420 01/12/01 07:48 meekins Target lifter vernier readings on jan 10 2001
  • 54419 01/12/01 06:41 Lingyan HV change for S1, gas and aerogel
  • 54416 01/12/01 05:56 kinney re: hallasc16
  • 54405 01/12/01 04:43 atrig Trigger-Download
  • 54404 01/12/01 04:39 atrig Trigger-Download
  • 54403 01/12/01 04:34 Nilanga Missing folis of the optics target
  • 54402 01/12/01 04:26 atrig Trigger-Download
  • 54392 01/12/01 03:19 Lingyan Left arm gas TDC back
  • 54389 01/12/01 02:34 dohrmann new calib for right arm shower and preshower
  • 54388 01/12/01 02:05 C. Williamso detector alignment
  • 54387 01/12/01 02:03 jpchen target postion and lifter
  • 54384 01/12/01 00:41 kinney Target/rate mystery
  • Thursday

  • 54369 01/11/01 22:55 markowitz magnet screen shot
  • 54361 01/11/01 21:32 Markowitz y_target
  • 54358 01/11/01 20:29 X. Zheng left arm gas cherenkov problem
  • 54357 01/11/01 20:17 X. Zheng Gas cherenkov calibration
  • 54352 01/11/01 19:28 Adaq spot
  • 54345 01/11/01 16:45 gilman hall a radiator position calibrated
  • 54339 01/11/01 14:35 Saha Arc Integral Measurement
  • 54336 01/11/01 14:21 gilman beam coming back
  • 54334 01/11/01 13:22 gilman beam soon
  • 54333 01/11/01 12:38 Reinhold contolled access
  • 54332 01/11/01 12:30 gilman down time ending, new radiator software
  • 54331 01/11/01 11:52 LeRose Beam Status Update
  • 54330 01/11/01 11:47 LeRose Beam Status
  • 54325 01/11/01 10:24 LeRose Right Hut AC
  • 54321 01/11/01 10:10 LeRose Checklist N2 levels
  • 54317 01/11/01 09:57 LeRose Spectrometer Move
  • 54316 01/11/01 09:38 gomez Re.: ioc hallasc16 crash
  • 54315 01/11/01 08:40 Calarco DAQ Deadtime
  • 54311 01/11/01 08:15 kinney A-Shift Summary
  • 54310 01/11/01 08:13 kinney Deadtime Understanding
  • 54306 01/11/01 07:59 wangxu attention
  • 54292 01/11/01 06:59 kinney RHS magnet ioc crash
  • 54282 01/11/01 04:58 Lingyan Software problem solved
  • 54277 01/11/01 04:33 wangxu Espace and Gas Cherenkov Status
  • 54263 01/11/01 00:07 Adaq shift summary
  • Wednesday

  • 54253 01/10/01 22:56 x. jiang pion rejector hv setting before run 1168
  • 54249 01/10/01 22:08 Johna HRS-Right PS and TS HV settings.
  • 54221 01/10/01 15:19 R. Michaels L-arm fastbus problems
  • 54217 01/10/01 14:24 Saha Alarm Handler, S1 PMT HV off
  • 54184 01/10/01 09:13 Lingyan AM HV change
  • 54181 01/10/01 07:55 kinney A-shift report
  • 54178 01/10/01 07:40 Lingyan A1 HV change
  • 54177 01/10/01 07:05 C. Williamso High Voltage Electron Arm
  • 54176 01/10/01 06:57 C. Williamso television camera Hall A
  • 54175 01/10/01 06:55 C. Williamso Q2 left HRS liquid nitrogen
  • 54174 01/10/01 06:04 kinney A-Dump water fault
  • 54173 01/10/01 03:58 kinney A-Dump water fault
  • 54172 01/10/01 03:01 kinney A-Dump water Fault
  • 54171 01/10/01 00:04 Markowitz shift summary
  • Tuesday

  • 54169 01/09/01 22:49 johna HRS-left Preshower HV changes.
  • 54166 01/09/01 22:02 Markowitz epics hv gui
  • 54163 01/09/01 20:27 Adaq faraday cup
  • 54162 01/09/01 20:26 K. McCormick HV ready and directions
  • 54152 01/09/01 18:43 R. Michaels fastbus busy hanging up
  • 54138 01/09/01 15:18 gilman radiator status
  • 54137 01/09/01 15:14 gilman fpp ch 3 and 4 turned off
  • 0 01/09/01 15:02 no_name no_key
  • 54120 01/09/01 12:43 Michaels/Cal Checks of Trigger
  • 54118 01/09/01 11:50 atrig Trigger-Download
  • 54116 01/09/01 11:45 atrig Trigger-Download
  • 54114 01/09/01 11:40 atrig Trigger-Download
  • 54112 01/09/01 11:36 atrig Trigger-Download
  • 54110 01/09/01 11:22 atrig Trigger-Download
  • 54108 01/09/01 11:14 atrig Trigger-Download
  • 54106 01/09/01 11:07 atrig Trigger-Download
  • 54104 01/09/01 10:57 atrig Trigger-Download
  • 54102 01/09/01 10:49 atrig Trigger-Download
  • 54100 01/09/01 10:42 atrig Trigger-Download
  • 54096 01/09/01 10:29 atrig Trigger-Download
  • 54054 01/09/01 05:58 atrig Trigger-Download
  • 54053 01/09/01 05:55 atrig Trigger-Download
  • Monday

  • 54049 01/08/01 22:13 atrig Trigger-Download
  • 54042 01/08/01 21:57 R. Michaels BCM L-arm restored, helicity bad
  • 54043 01/08/01 21:57 atrig Trigger-Download
  • 54040 01/08/01 19:05 C. Williamso A1-A2 position alignment
  • 54023 01/08/01 12:49 atrig Trigger-Download
  • 54022 01/08/01 12:29 atrig Trigger-Download
  • 54021 01/08/01 12:27 atrig Trigger-Download
  • 54019 01/08/01 12:22 atrig Trigger-Download
  • 54018 01/08/01 12:19 atrig Trigger-Download
  • 54017 01/08/01 12:18 atrig Trigger-Download
  • 54015 01/08/01 12:10 atrig Trigger-Download
  • 54014 01/08/01 12:03 atrig Trigger-Download
  • 54013 01/08/01 11:55 atrig Trigger-Download
  • 54012 01/08/01 11:53 atrig Trigger-Download
  • 54011 01/08/01 11:50 atrig Trigger-Download
  • 54010 01/08/01 10:43 atrig Trigger-Download
  • Sunday

  • 53999 01/07/01 16:40 atrig Trigger-Download
  • 53998 01/07/01 16:40 atrig Trigger-Download
  • 53997 01/07/01 16:38 atrig Trigger-Download
  • 53996 01/07/01 16:37 atrig Trigger-Download
  • Saturday

  • 53982 01/06/01 15:50 atrig Trigger-Download
  • 53979 01/06/01 14:43 atrig Trigger-Download
  • 53978 01/06/01 14:42 atrig Trigger-Download
  • 53977 01/06/01 14:41 atrig Trigger-Download
  • 53974 01/06/01 14:13 atrig Trigger-Download
  • 53972 01/06/01 14:08 R. Michaels stuff unplugged
  • Friday

    
    

    Thursday

  • 53942 01/04/01 21:21 R. Michaels DAQ status
  • Wednesday

  • 53924 01/03/01 18:09 R. Michaels more modifications
  • 53923 01/03/01 15:26 R. Michaels tests of trigger download - ignore
  • 53922 01/03/01 15:10 atrig Trigger-Download
  • 53921 01/03/01 15:06 atrig Download
  • 53920 01/03/01 15:05 atrig Download
  • 53919 01/03/01 15:01 atrig tst_map
  • 53919 01/03/01 14:58 atrig tst_map
  • Tuesday

  • 53914 01/02/01 17:22 R. Michaels DAQ L-arm config, detmap mods